OpenStudy (anonymous):

Solve the following proportion:

OpenStudy (anonymous):

\[\frac{ x }{ 4 }=\frac{ 7 }{ 14 }\]

OpenStudy (anonymous):

@hartnn , Do I just cross multiply?

OpenStudy (anonymous):

@campbell_st @hero

hero (hero):

Multiply both sides by 4, then simplify the right side.

OpenStudy (anonymous):

\[\frac{ 4x }{ 16 }=\frac{ 1 }{ 2 }\]

OpenStudy (anonymous):

You multiplied by 4/4 Just multiply by 4 on both sides and simplify.

hero (hero):

Actually, reducing the right side was a good start, but you should have ended up with \[\frac{x}{4} = \frac{1}{2}\] Now you still have to multiply both sides by 4 to isolate x.

OpenStudy (anonymous):

I'm so lost...

OpenStudy (anonymous):

\[\frac{x}{a} = \frac{b}{c}\] Multiply both sides by a: \[a \frac{x}{a} = a \frac{b}{c}\] The a's on the left side cancel: \[x = a \frac{b}{c}\]

OpenStudy (anonymous):

\[x=\frac{ 1 }{ 2 }(4)\]

OpenStudy (anonymous):

Yep. Which simplifies to what?

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

There you go :)